OpenStudy (anonymous):

Solve for x: -4(x - 3) = 16 x = 1 x = -1 x = 7 x = -7

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

-4x + 12 = 16 -4x = 4 x = -1